Find the first term and the common difference of the arithmetic sequence whose 5th term is 7 and 13th term is 39.
Answer by richwmiller(17219)

an = a + (n - 1)d
39 = a + 12d
7 = a + 4d,
32=8d
4=d
39 = a + 48
-9=a1